# Glitch in Time: Exploiting Temporal Misalignment of IMU For Eavesdropping

### v2

@@ 6. Discussion
Countermeasures: To prevent the STAG-based zero-permission attack for unauthorized speech recognition, several countermeasures can be implemented. First, imposing stricter access controls on motion sensor data can significantly reduce the risk of exploitation. This includes requiring explicit user permissions for accessing accelerometer and gyroscope data, even at lower sampling rates. This measure ensures that only authorized applications can access sensitive sensor data, thereby protecting user privacy.
Another effective measure would be to employ noise injection techniques, which add benign noise to sensor readings. This approach obfuscates the data, rendering it less useful for precise speech recognition without significantly degrading the performance of legitimate applications that rely on sensor data.
Finally, configuring the magnetometer and IMU to be directly connected to the host, rather than as a slave, would help mitigate the issue. In this setup, the magnetometer interrupts would not affect the FIFO buffer within the IMU, as they would be handled by the host instead.
By addressing these potential vulnerabilities and implementing robust countermeasures, the risk of unauthorized speech recognition via STAG can be significantly mitigated. These measures ensure better protection of user privacy while maintaining the functionality and performance of authorized applications.

### v3

@@ 6. Discussion
Countermeasures: Several effective countermeasures can be implemented to mitigate STAG-based attacks that exploit zero-permission access to IMUs for eavesdropping. These are summarized in Table 7.
  • Stricter Access Controls: Enforcing stricter access controls on motion sensor data is crucial. Requiring explicit user permissions for accessing accelerometer and gyroscope data, even at lower sampling rates, ensures that only authorized applications can access these sensitive sensors. This significantly reduces the risk of covert speech recognition.
  • Noise Injection: Introducing noise injection techniques can obfuscate sensor data, hindering attackers’ ability to reconstruct accurate speech patterns. By adding benign noise to sensor readings, legitimate applications can still use the data while degrading its usefulness for malicious purposes. This approach balances user privacy with the functionality of authorized applications.
  • Hardware Modifications: Modifying hardware configurations by connecting the magnetometer and IMU directly to the host system rather than as peripheral devices enhance security. This configuration prevents magnetometer interrupts from affecting the IMU’s data buffer (a temporary storage area), thereby isolating sensor data streams and eliminating potential channels for unauthorized data access.
Implementing these countermeasures significantly diminishes the risk of eavesdropping via STAG-based attacks. This comprehensive strategy provides robust user privacy protection while maintaining the functionality of legitimate, sensor-dependent applications.